# A Survey on Diffusion Models for Recommender Systems

### v1

@@ 2.2.2. Score-based Generative Models (SGMs)
Score-based generative models (SGMs) (Song et al., 2020) further generalize DDPM’s discrete diffusion processes to a continuous framework based on stochastic differential equations (SDEs). For clarity, here we adopt the notation $t\in[0,T]$ for SGMs instead of the step size $k=1,\dots,K$ in DDPMs. Consequently, the sequence $\mathbf{x}_{0},\dots,\mathbf{x}_{K}$ is replaced with a continuous function $\mathbf{x}(t)$.
@@
Upon finishing the training process, we can generate samples based on various techniques like the Euler-Maruyama (EM) (Mao, 2015), Prediction-Correction (PC), or Probability Flow ODE method (Song et al., 2020).
2.2.3. Diffusion Model Guidance
@@
Although classifier guidance is a common and versatile approach to improve the sample quality, it heavily relies on the availability of a noise-robust pre-trained classifier $p_{\phi}(c|\mathbf{x})$. This requirement largely depends on the existence of annotated data to well train the classifier network, which is impractical in many real-world data-hungry applications. To this end, the classifier-free guidance is proposed. Compared to the high accuracy of the labeled conditional diffusion model, the sampling under unlabeled conditions solely relies on self-information for guidance, and is better at generating innovative, creative and diverse data samples (Choi et al., 2021; Epstein et al., 2023; Chao et al., 2022; Kollovieh et al., 2024).
As a result, the conditional diffusion models have been widely adopted in various online application services due to their high-quality and well-controlled generative outputs (Bansal et al., 2023; Zhang et al., 2023a; Kollovieh et al., 2024; Nichol et al., 2022).
@@ 3. Diffusion Models for Recommendation: Taxonomy
Based on the decomposition of modern recommender systems discussed in Section 2.1, we introduce the taxonomy framework of diffusion models for recommendation according to different roles that diffusion models play at different parts of the modern deep learning based recommender system pipeline: (1) diffusion for data engineering & encoding, (2) diffusion as recommender model, and (3) diffusion for content presentation. The overall taxonomy framework is depicted in Figure 3.
Figure 3. The overall categorization of diffusion models for recommendation.
@@ 3.1. Diffusion for Data Engineering & Encoding
Data engineering and encoding encompass the sophisticated processes of refining and converting the vast array of raw data gathered from online sources into structured formats or advanced neural embeddings, tailored for the efficient functioning of downstream recommendation systems. While the current recommender systems generally suffer from problems like data sparsity (Wu et al., 2023) and noisy data (Zhao et al., 2024a), diffusion models emerge as an exceptionally potent category of generative models to mitigate such issues during the data engineering and encoding stage. They exhibit outstanding prowess in the dual realms of data augmentation and representation enhancement, which are pivotal in significantly bolstering the overall effectiveness and precision of downstream recommendation performance. Diffusion for data augmentation refers to the process of manipulating and augmenting the original raw training data (e.g., ID features, user behavior sequences, and user-item interactions), while diffusion for feature enhancement primarily focuses on strengthening and enhancing the neural embeddings (e.g., user & item embeddings, and multi-modal representations) during the data encoding phase for downstream recommendation models.
@@ 3.1.1. Diffusion for Data Augmentation
As a powerful class of generative models, diffusion models are able to capture the underlying distribution of given data and synthesize realistic samples for the downstream recommenders. By augmenting the original training data, diffusion models can thereby improve the recommender systems from various perspectives, e.g., handling sparsity, enhancing diversity, reducing noise, and improving model generalization. In the following, we will discuss leveraging diffusion models for data augmentation according to different data modalities to be generated for augmentation, i.e., sequential data augmentation, feature imputation, user-item interaction synthesis, and side information editing.
@@
Apart from the sequential data, multi-field categorical/numerical features (or say tabular data) are also of great importance for recommender systems to capture the dynamic user preferences by modeling feature crossings (Lin et al., 2023a; Guo et al., 2017; Wang et al., 2017). Hence, diffusion models are widely utilized for feature generation and imputation (Villaizán-Vallelado et al., 2024; Kotelnikov et al., 2023; Zhang et al., 2023b; Lee et al., 2023). For example, TabDDPM introduces the design of DDPM (Ho et al., 2020) to the field of tabular data, and enables it to synthesizing a mixed types of data including ordinal, numerical, and categorical features. TabSyn (Zhang et al., 2023b) involves the diffusion model within a variational autoencoder (VAE) crafted latent space, and achieves high-quality for mixed-type feature generation with faster generation speed via a linear noise schedule (i.e., less than 20 reverse steps). Villaizán-Vallelado et al. (2024) further incorporate a full encoder-decoder transformer with diffusion process as the denoising model, which allows for the conditioning attention mechanism while effectively capturing and representing complex interactions and dependencies among the input features.
Other works leverage diffusion models to simulate the user-item interactions as the augmented training data for recommendation models with various purposes, e.g., cold-start problems (Wu et al., 2023), privacy-preserving concerns (Lilienthal et al., 2024), and adversarial robustness (Liu et al., 2024c). For instance, Liu et al. (2024c) propose a target-oriented diffusion attack model to generate deceptive user interactions profiles with guidance of cross-attention mechanisms for shilling attacks towards the target item, which helps deepen the insights of the vulnerabilities and adversarial attacks in recommender systems. SDRM (Lilienthal et al., 2024) employs diffusion models to capture complex patterns of real-world datasets, and generates high-quality synthetic user-item interactions to augment or even replace the original dataset, aiming at addressing the data sparsity problem, as well as the privacy and security concerns for recommendation.
@@ 3.1.2. Diffusion for Representation Enhancement
Different from directly augmenting the raw input data during the data engineering phase as introduced above, diffusion for representation enhancement generally focuses on capturing the underlying distribution of raw input data and transforming it into dynamic and robust feature embeddings to assist the training of downstream recommendation models. As a unique instance of self-supervised learning paradigms with explicit denoising process, diffusion models are capable of establishing generalized latent spaces for enhanced representations and therefore addressing the key challenges of recommender systems, e.g., the multi-interest and ever-evolving user preference (Liu et al., 2024b; Xi et al., 2024), the multiple latent aspects of items (Fan et al., 2021, 2022), and the noise and uncertainty of interaction data (Hurley and Zhang, 2011; Wang et al., 2021b). A range of related studies (Zhao et al., 2024a; Yi et al., 2024; He et al., 2024b; Zhao et al., 2024b; Ma et al., 2024d; Wang et al., 2024b; Zuo and Zhang, 2024; Li et al., 2024e) have been proposed in this line of research, each of which incorporates diffusion models for enhanced representation learning under different goals and scenarios. We briefly introduce these research works as follows.
@@
DDRM (Zhao et al., 2024a), DiffGT (Yi et al., 2024) and MISD (Li et al., 2024e) aim to denoise the implicit user feedback. DDRM (Zhao et al., 2024a) attempts to robustify the user and item representations for arbitrary recommendation models. It injects controlled Gaussian noise into user and item embeddings during a forward phase and then iteratively removes this noise during a reverse denoising phase, guided by a specialized denoising module utilizing the collaborative signals. Besides, in the inference stage, DDRM leverages the average embeddings of the user’s historically interacted items as the starting point rather than a pure noise vector to further promote the personalization. DiffGT (Yi et al., 2024) further considers handling the noisy implicit user feedback for neural graph recommenders. The authors showcase the anisotropic nature of recommendation data and propose to incorporates anisotropic directional Gaussian noise to improve the diffusion process, ensuring that the forward noise better aligns with the observed characteristics of recommendation data. The model also integrates a graph transformer architecture with a linear attention module to efficiently robustify the noisy embeddings, which is guided by personalized information to improve the estimation of user preferences. MISD (Li et al., 2024e), on the other hand, leverages diffusion models to address the noisy feedback during multi-interest user modeling for multi-behavior sequential recommendation.
@@
CausalDiffRec (Zhao et al., 2024b) employs causal diffusion models to address the out-of-distribution (OOD) data in the field of graph-enhanced recommendation. It incorporates backdoor adjustment and variational inference to capture the real environmental distribution, and then uses it as prior knowledge to guide the reverse phase of the diffusion process for invariant representation learning, which eliminates the impact of environmental confounders. The authors also provide theoretical derivations to prove that the proposed objective of CausalDiffRec encourages the model to learn environment-invariant graph representations, achieving excellent generalization performance in recommendations under distribution shifts and OOD data.
@@
RDRec (He et al., 2024b) focuses on improving the review-based textual embeddings to better model the diverse and dynamic user interests when facing different items. The model corrupts the user representations by adding noises to the original review-based textual features of the user interaction sequence, and the perturbed user representations undergoes denoising via a transformer approximator with the awareness of target item information. Consequently, the model learns to capture the dynamic user preferences and produce generalized user interest embeddings for final recommendation.
@@
Diff-MSR (Wang et al., 2024b) concentrates on the cold-start problem in multi-scenario recommendation, and utilize the transfer capabilities of diffusion models to enhance the representation learning of long-tail cold-start scenarios with the help of other scenarios with sufficient training data. Built upon a pretrained multi-scenario recommendation model, the authors design a piece-wise variance schedule, and then train a cold-or-rich domain classifier to obtain the candidates from rich domains (if incorrectly classified) to generate high-quality and informative embeddings for cold-start domains. It turns out that diffusion models are capable of capturing the commonality and distinction of various scenarios, enabling effective knowledge transferring among cold-start and rich domains.
@@ 3.1.3. Discussion
When adapting diffusion models for the data engineering & encoding stage, the diffusion models have showcased various impressive characteristics, e.g., powerful and controllable generative capability, high-quality and diverse output, robust latent space representation, and high flexibility. Among them, we argue that the flexibility serves as the core characteristic, where diffusion models generally play the role of a bridge and a connector, being compatible with different upstream input data types (e.g., texts, images, and user-item interactions) and different downstream recommendation models (e.g., collaborative recommenders, sequential recommenders, and graph-enhanced recommenders). This is attributed to the fact that diffusion models are actually a general self-supervised learning paradigm with the fundamental design principle of gradually transforming noise into structured data through an iterative denoising process, which makes the diffusion model a versatile plug-in toolkit for recommender systems.
3.2. Diffusion as Recommender Model
@@
The recommender model serves as the core component of the recommender system pipeline to select or rank the top-relevant items to satisfy users’ information needs based on the outcome of the previous data engineering & encoding stage. When adapting diffusion models as recommender models, the input for this stage can be structured data (e.g., user behavior sequence), neural embeddings from other encoders, or a combination of both, depending on the architecture design of the diffusion-based recommender. In this section, as shown in Figure 3, we mainly focus on the input formats and task formulations of diffusion-based recommender models, and thereby classify related research works into three categories: (1) collaborative recommendation, (2) context-aware recommendation, and (3) other applications.
@@
In general, collaborative recommendation, also known as collaborative filtering, is the basic recommendation task of modeling user preference primarily based on the user-item interaction records (i.e., user-item co-occurrence matrix). It is simple, straightforward, yet effective to capture the overall user preferences via the user behavior history. Built upon the basic collaborative signals, context-aware recommendation further considers the context information to enable more dynamic and more accurate user interest modelings. In this paper, the definition scope of context information is relatively broad, including but not limited to circumstances (e.g., time, location, and device), user temporal dynamic (e.g., behavior sequences instead of a set of records), and item attributes (e.g., textual descriptions and thumbnails). Lastly, other applications is referred to the tasks that are closely related to the recommendation scenarios but somehow differ from the aforementioned two categories in problem formulations and solution paradigms, such as cross-domain recommendation, learning to ranking, and computational advertising.
@@
Before elaborating on the details of this section, we would like to further clarify the key difference between “diffusion as recommender model” and “diffusion for data engineering & encoding”. As discussed in Section 3.1.3, the methods in the category of “diffusion for data engineering & encoding” focus on producing enhanced structured data or neural embeddings for other downstream recommender models. However, in this section, the diffusion models are directly trained as the recommender models to generate the user preference distribution over the item space, or produce latent user interest representations for item matching.

### v2

@@ 2.2.2. Score-based Generative Models (SGMs)
Score-based generative models (SGMs) (Song and Ermon, 2019; Song et al., 2020) further generalize DDPM’s discrete diffusion processes to a continuous framework based on stochastic differential equations (SDEs) (Van Kampen, 1976). For clarity, here we adopt the notation $t\in[0,T]$ for SGMs instead of the step size $k=1,\dots,K$ in DDPMs. Consequently, the sequence $\mathbf{x}_{0},\dots,\mathbf{x}_{K}$ is replaced with a continuous function $\mathbf{x}(t)$. Note that here we discuss the continuous property of the diffusion-denoising process, while the data itself can be either continuous (e.g., images) or discrete (e.g., texts).
@@
Upon finishing the training process, we can generate samples based on various techniques like the Euler-Maruyama (EM) (Mao, 2015), Prediction-Correction (PC) (Butcher, 2016), or Probability Flow ODE method (Song et al., 2020).
2.2.3. Conditional Diffusion Models
@@
Although classifier guidance is a common and versatile approach to improve the sample quality, it heavily relies on the availability of a noise-robust pre-trained classifier $p_{\phi}(c|\mathbf{x})$. This requirement largely depends on the existence of annotated data to well train the classifier network, which is impractical in many real-world data-hungry applications. To this end, the classifier-free guidance is proposed. Compared to the high accuracy of the labeled conditional diffusion model, the sampling under unlabeled conditions solely relies on self-information for guidance, and is better at generating innovative, creative and diverse data samples (Choi et al., 2021; Epstein et al., 2023; Chao et al., 2022; Kollovieh et al., 2024). The classifier-free guidance is usually the dense representations produced by specific encoder models based on certain contextual information like textual data, sequential data and graphic data.
As a result, both the classifier-based and classifier-free conditional diffusion models have been widely adopted in various online application services due to their high-quality and well-controlled generative outputs (Bansal et al., 2023; Zhang et al., 2023a; Kollovieh et al., 2024; Nichol et al., 2022).
@@ 3. Diffusion Models for Recommendation: Taxonomy
Based on the decomposition of modern recommender systems discussed in Section 2.1, we introduce the taxonomy framework of diffusion models for recommendation according to different roles that diffusion models play at different parts of the modern deep learning based recommender system pipeline: (1) diffusion for data engineering & encoding, (2) diffusion as recommendation model, and (3) diffusion for content presentation. We depict the overall taxonomy framework in Figure 3, and list the research works of each category to facilitate readers in quickly indexing and finding the corresponding works.
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=black, rounded corners, minimum width=2.5em, s sep=6pt, inner xsep=3pt, inner ysep=5pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=5.6em,font=, where level=2text width=5.3em,font=, where level=3text width=9.7em,font=, [Diffusion Models for
Recommendation, level_zero, text width=8.5em [Diffusion for Data
Engineering & Encoding, level_one [Data Augmentation,level_two [Sequential Data: PDRec (Ma et al., 2024b), Diff4Rec (Wu et al., 2023), CaDiRec (Cui et al., 2024), DiffuASR (Liu et al., 2023b), DGFedRS (Di et al., [n. d.]), SeeDRec (Ma et al., 2024c),level_three] [Multi-field Features: TabDDPM (Kotelnikov et al., 2023), TabSyn (Zhang et al., 2023b), MTabGen (Villaizán-Vallelado et al., 2024), CoDi (Lee et al., 2023),level_three] [User-item Interactions: ToDA (Liu et al., 2024c), SDRM (Lilienthal et al., 2024), SCONE (Lee et al., 2024),level_three] [Side Information: IPDGI (Chen et al., 2023), DiffKG (Jiang et al., 2024b), DICES (Dong et al., 2024), DiffMM (Jiang et al., 2024a), SimDiff (Li et al., 2024b),level_three] ] [Representation Enhancement,level_two [DDRM (Zhao et al., 2024a), DiffGT (Yi et al., 2024), MISD (Li et al., 2024e), Diff-DGMN (Zuo and Zhang, 2024), CausalDiffRec (Zhao et al., 2024b), RDRec (He et al., 2024b), MCDRec (Ma et al., 2024d), Diff-MSR (Wang et al., 2024b),level_three] ] ] [Diffusion as
Recommendation Model, level_one [Collaborative Recommendation,level_two [BSPM (Choi et al., 2023), DiffRec (Wang et al., 2023b), CODIGEM (Walker et al., 2022), CFDiff (Hou et al., 2024), DGRM (Jiangzhou et al., 2024), RecFusion (Bénédict et al., 2023), GiffCF (Zhu et al., 2024), DIEXRS (Guo et al., 2023), FairDR (Malitesta et al., 2024),level_three] ] [Context-aware Recommendation,level_two [DreamRec (Yang et al., 2024c), LeadRec (Wang et al., 2024d), DiffRIS (Niu et al., 2024), DimeRec (Li et al., 2024c), CDDRec (Wang et al., 2024c), DiffRec (Du et al., 2023), DiffuRec (Li et al., 2023a), RecDiff (Li et al., 2024f), DSDRec (Wang et al., 2024e), Diff-POI (Qin et al., 2023), DCPR (Long et al., 2024), LD4MRec (Yu et al., 2023a), CGSoRec (He et al., 2024a),level_three] ] [Other Applications,level_two [DiffCDR (Xuan, 2024), DMSR (Tomasi et al., 2024), DCDR (Lin et al., 2024a), DiffBid (Guo et al., 2024), ,level_three] ] ] [Diffusion for
Content Presentation, level_one [Fashion Generation,level_two [DiFashion (Xu et al., 2024a),level_three] ] [Ad Creative Generation,level_two [AdBooster (Shilova et al., 2023), CG4CTR (Yang et al., 2024d), DynaPIG (Czapp et al., 2024),level_three] ] [General Content Generation,level_two [MMCRec (Mukande et al., 2024), PMG (Shen et al., 2024), GeneRec (Wang et al., 2023a),level_three] ] ] ] ]
Figure 3. The overall categorization of diffusion models for recommendation. We also list the research works of each category attached with the corresponding method name and reference.
@@ 3.1. Diffusion for Data Engineering & Encoding
Data engineering and encoding encompass the sophisticated processes of refining and converting the vast array of raw data gathered from online sources into structured formats or advanced neural embeddings, tailored for the efficient functioning of downstream recommendation systems. While the current recommender systems generally suffer from problems like data sparsity (Wu et al., 2023), noisy data (Zhao et al., 2024a) and filtering bubble (Piao et al., 2023), diffusion models emerge as an exceptionally potent category of generative models to mitigate such issues during the data engineering and encoding stage. They exhibit outstanding prowess in the dual realms of data augmentation and representation enhancement, which are pivotal in significantly bolstering the overall effectiveness and precision of downstream recommendation performance. Diffusion for data augmentation refers to the process of manipulating and augmenting the original raw training data (e.g., ID features, user behavior sequences, and user-item interactions) in data engineering phase, while diffusion for representation enhancement primarily focuses on strengthening and enhancing the neural embeddings (e.g., user & item embeddings, and multi-modal representations) during the data encoding phase for downstream recommendation models.
@@ 3.1.1. Diffusion for Data Augmentation
As a powerful class of generative models, diffusion models are able to capture the underlying distribution of given data and synthesize realistic samples for the downstream recommenders. By augmenting the original training data, diffusion models can thereby improve the recommender systems from various perspectives, e.g., handling sparsity, enhancing diversity, reducing noise, and improving model generalization. In the following, we will discuss leveraging diffusion models for data augmentation according to different data types to be generated for augmentation, i.e., sequential data augmentation, feature imputation, user-item interaction synthesis, and side information editing.
@@
Apart from the sequential data, multi-field categorical/numerical features (or say tabular data) are also of great importance for recommender systems to capture the dynamic user preferences by modeling feature crossings (Lin et al., 2023a; Guo et al., 2017; Wang et al., 2017). Hence, diffusion models are widely utilized for feature generation and imputation (Villaizán-Vallelado et al., 2024; Kotelnikov et al., 2023; Zhang et al., 2023b; Lee et al., 2023). For example, TabDDPM (Kotelnikov et al., 2023) introduces the design of DDPM (Ho et al., 2020) to the field of tabular data, and enables it to synthesizing a mixed types of data including ordinal, numerical, and categorical features. TabSyn (Zhang et al., 2023b) involves the diffusion model within a variational autoencoder (VAE) crafted latent space, and achieves high-quality for mixed-type feature generation with faster generation speed via a linear noise schedule (i.e., less than 20 reverse steps). Villaizán-Vallelado et al. (2024) further incorporate a full encoder-decoder transformer with diffusion process as the denoising model, which allows for the conditioning attention mechanism while effectively capturing and representing complex interactions and dependencies among the input features.
Other works leverage diffusion models to simulate the user-item interactions as the augmented training data for recommendation models with various purposes, e.g., cold-start problems (Wu et al., 2023), hard negative sample mining (Lee et al., 2024), privacy-preserving concerns (Lilienthal et al., 2024), and adversarial robustness (Liu et al., 2024c). For instance, Liu et al. (2024c) propose a target-oriented diffusion attack model to generate deceptive user interactions profiles with guidance of cross-attention mechanisms for shilling attacks towards the target item, which helps deepen the insights of the vulnerabilities and adversarial attacks in recommender systems. SDRM (Lilienthal et al., 2024) employs diffusion models to capture complex patterns of real-world datasets, and generates high-quality synthetic user-item interactions to augment or even replace the original dataset, aiming at addressing the data sparsity problem, as well as the privacy and security concerns for recommendation.
@@ 3.1.2. Diffusion for Representation Enhancement
Different from directly augmenting the raw input data during the data engineering phase as introduced above, diffusion for representation enhancement generally focuses on capturing the underlying distribution of raw input data and transforming it into robust feature embeddings to assist the training of downstream recommendation models. As a unique instance of self-supervised learning paradigms with explicit denoising process, diffusion models are capable of establishing generalized latent spaces for enhanced representations and therefore addressing several key challenges of recommender systems, e.g., the multi-interest and ever-evolving user preference (Liu et al., 2024b; Xi et al., 2024), the diverse item aspects (Fan et al., 2021, 2022a), and the noise and uncertainty of interaction data (Hurley and Zhang, 2011; Wang et al., 2021b). A range of related studies (Zhao et al., 2024a; Yi et al., 2024; He et al., 2024b; Zhao et al., 2024b; Ma et al., 2024d; Wang et al., 2024b; Zuo and Zhang, 2024; Li et al., 2024e) have been proposed in this line of research, each of which incorporates diffusion models for enhanced representation learning under different goals and scenarios. We briefly introduce these research works as follows.
@@
DDRM (Zhao et al., 2024a), DiffGT (Yi et al., 2024) and MISD (Li et al., 2024e) aim to denoise the implicit user feedback. DDRM (Zhao et al., 2024a) attempts to robustify the user and item representations for arbitrary recommendation models. It injects controlled Gaussian noise into user and item embeddings during a forward phase and then iteratively removes this noise during a reverse denoising phase, guided by a specialized denoising module utilizing the collaborative signals. Besides, in the inference stage, DDRM leverages the average embeddings of the user’s historically interacted items as the starting point rather than a pure noise vector to further promote the personalization. DiffGT (Yi et al., 2024) further considers handling the noisy implicit user feedback for neural graph recommenders. The authors showcase the anisotropic nature of recommendation data and propose to incorporates anisotropic directional Gaussian noise to improve the diffusion process, ensuring that the forward noise better aligns with the observed characteristics of recommendation data. The model also integrates a graph transformer architecture with a linear attention module to efficiently robustify the noisy embeddings, which is guided by personalized information to improve the estimation of user preferences. MISD (Li et al., 2024e), on the other hand, leverages diffusion models to address the noisy feedback in multi-interest user modeling, targeting multi-behavior sequential recommendation.
@@
CausalDiffRec (Zhao et al., 2024b) employs causal diffusion models to address the out-of-distribution (OOD) issue in the field of graph-enhanced recommendation. It incorporates backdoor adjustment and variational inference to capture the real environmental distribution, and then uses it as prior knowledge to guide the reverse phase of the diffusion process for invariant representation learning, which eliminates the impact of environmental confounders. The authors also provide theoretical derivations to prove that the proposed objective of CausalDiffRec encourages the model to learn environment-invariant graph representations, achieving excellent generalization performance in recommendations under distribution shifts and OOD data.
@@
RDRec (He et al., 2024b) focuses on improving the review-based textual embeddings to better model the diverse and dynamic user interests when facing different items. The model corrupts the user representations by adding noises to the original review-based textual features of the user interaction sequence, and the perturbed user representations undergo denoising via a transformer approximator with the awareness of target item information. Consequently, the model is able to learn to capture the dynamic user preferences and therefore generate generalized user interest embeddings for final recommendation.
@@
Diff-MSR (Wang et al., 2024b) concentrates on the cold-start problem in multi-scenario recommendation, and leverages the transfer capabilities of diffusion models to enhance the representation learning of long-tail cold-start scenarios, utilizing information from other scenarios with ample training data. Built upon a pretrained multi-scenario recommendation model, the authors design a piece-wise variance schedule, and then train a cold-or-rich domain classifier to obtain the candidates from rich domains (if incorrectly classified) to generate high-quality and informative embeddings for cold-start domains. It turns out that diffusion models are capable of capturing the commonality and distinction of various scenarios, enabling effective knowledge transferring among cold-start and rich domains.
@@ 3.1.3. Discussion
When adapting diffusion models for the data engineering & encoding stage, the diffusion models have showcased various impressive characteristics, e.g., powerful and controllable generative capability, high-quality and diverse output, robust latent space representation, and high flexibility. Among them, as illustrated in Figure 4, we argue that the flexibility serves as the core characteristic, where diffusion models generally play the role of a bridge and a connector, being compatible with different upstream input data types (e.g., tabular data, user behavior sequence, multi-modal data, and graph data) and different downstream recommendation models (e.g., general recommenders, sequential recommenders, multi-modal recommenders, and graph-enhanced recommenders). This is attributed to the fact that diffusion models are actually a general self-supervised learning paradigm with the fundamental design principle of gradually transforming noise into structured data through an iterative denoising process, which makes the diffusion model a versatile plug-in toolkit for recommender systems.
3.2. Diffusion as Recommendation Model
@@
The recommendation model serves as the core component of the recommender system pipeline to select or rank the top-relevant items to satisfy users’ information needs based on the outcome of the previous data engineering & encoding stage. When adapting diffusion models as recommendation models, the input for this stage can be structured data (e.g., user behavior sequence), neural embeddings from other encoders, or a combination of both, depending on the architecture design of the diffusion-based recommender. In this section, as shown in Figure 3, we mainly focus on the input formats and task formulations of diffusion-based recommendation models, and thereby classify related research works into three categories: (1) collaborative recommendation, (2) context-aware recommendation, and (3) other applications.
@@
In general, collaborative recommendation, also known as collaborative filtering, is the basic recommendation task of modeling user preference primarily based on the user-item interaction records (i.e., user-item co-occurrence matrix). It is simple, straightforward, yet effective to capture the overall user preferences via the user-item interactions. Built upon the basic collaborative signals, context-aware recommendation further considers the context information to enable more dynamic and more accurate user interest modelings. In this paper, the definition scope of context information is relatively broad, including but not limited to circumstances (e.g., time, location, and device), user temporal dynamic (e.g., behavior sequences instead of a set of records), and item attributes (e.g., textual descriptions and thumbnails). Lastly, other applications is referred to the tasks that are closely related to the recommendation scenarios but somehow differ from the aforementioned two categories in problem formulations and solution paradigms, such as cross-domain recommendation, learning to rank, and computational advertising.
@@
Before elaborating on the details of this section, we would like to further clarify the key difference between “diffusion as recommendation model” and “diffusion for data engineering & encoding”. As discussed in Section 3.1.3, the methods in the category of “diffusion for data engineering & encoding” focus on producing enhanced structured data or neural embeddings for other downstream recommendation models. However, in this section, the diffusion models are directly trained as the recommendation models to generate the user preference distribution over the item space, or produce latent user interest representations for item matching.